OpenStudy (anonymous):

2. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

3. Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

@alexwright think you help with this?

OpenStudy (anonymous):

hold on I can try

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

#2 [d = 0.5g t^2\] divide both sides by 0.5t^2\[\frac{ 2d }{ t^2 } = g\] they want the value of g if t = 10 and d = 490 \[g = \frac{ 2(490) }{ 10^2 } = 9.8\] the units is feet/s^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

#3 divide both sides by w2 \[d_2 = \frac{ w_1 \times d_1 }{ w_2 }\]

OpenStudy (anonymous):

which sides? the fractions sides or both sides of the = sign?

OpenStudy (anonymous):

@alexwright

OpenStudy (anonymous):

both

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so 0.5d=\[\frac{ 0.5w \times 0.5d }{ d }\]

OpenStudy (anonymous):

@alexwright right?

OpenStudy (anonymous):

yeah I think so

OpenStudy (anonymous):

ok thanks for the help on all the questions i asked

OpenStudy (anonymous):

@alexwright

OpenStudy (anonymous):

lol np thank you for your help too :)

OpenStudy (anonymous):

no problem